OpenStudy (anonymous):

the ratio of the measurement of an angle to it's complement is 3:2.find the measure of the angle

OpenStudy (anonymous):

let the angle be x, then its compliment is 90-x given : \[\frac{ x }{ 90-x }=\frac{ 3 }{2 }\]

OpenStudy (anonymous):

that's it already dhashni??

OpenStudy (anonymous):

no! u got to solve it for x ..... x is the measure of the angle

OpenStudy (anonymous):

by cross multiplying you get 2x=270-3x => 5x=270 =>x=54 the answer is 54

OpenStudy (anonymous):

thank you very much!!